A car of mass 9500kg is moving up a cliff from rest with a velocity of v=(3t^2 + 2t)î +5j m/s for 1 minute.If the slope of the cliff is 1,what distance was covered by the car in 1minute


1
Expert's answer
2021-01-28T17:09:34-0500
"x(t)=3\frac{t^3}{3}+2\frac{t^2}{2}=t^3+t^2\\y(t)=5t"

"x(60)=60^3+60^2=219600\ m\\y(60)=5(60)=300\ m\\s=\sqrt{219600^2+300^2}=219600\ m"